# Do AI Models Perform Human-like Abstract Reasoning Across Modalities?

### v1

@@ 6 Limitations
Ethics Statement
This paper uses data from a set of human studies described in Moskvichev et al. (2023). The authors of that study obtained an IRB exemption from the University of New Mexico IRB to perform their study. The data we obtained contained no identifying or other private information about the study participants. We did not identify any other ethics issues with the studies reported here.
Reproducibility Statement
Upon publication we will publish a web page for this paper with all data and code. The ConceptARC dataset is already publicly available at the website https://github.com/victorvikram/ConceptARC. The only reasons our results may not be fully reproducible are the non-deterministic nature of the AI models we evaluated, especially due to the fact that our reasoning models were restricted to Temperature 1, and the unpredictability of model releases and deprecations by the companies that created the proprietary models we used. In addition to derived rules and output grids, we collected all reasoning traces by different grids, either directly, or with the most detailed summary setting (OpenAI does not provide reasoning tokens directly). Further, we collected all Python calls by the models, which we interleaved with reasoning data in chronological order, to ensure that model outputs are fully documented.
@@ Acknowledgments
This work has been supported by the BANYAN project at Sandia National Laboratories and by the Templeton World Charity Foundation, Inc. (funder DOI 501100011730) under the grant https://doi.org/10.54224/20650. We thank Kaleda K. Denton for helpful comments on the manuscript.

### v2

@@ Acknowledgments
Sandia National Laboratories is a multimission laboratory managed and operated by National Technology and Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International, Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA-0003525. This work was conducted as part of the BANYAN Institute, funded by Sandia National Laboratories’ lab-directed research and development program. This work was also supported by the Templeton World Charity Foundation, Inc. (funder DOI 501100011730) under the grant https://doi.org/10.54224/20650. We thank Kaleda K. Denton for helpful comments on the manuscript.